In a sample of the non-stoichiometric crystal Fe0.96O,the % of cationic sites occupied by Fe3+ ions is?


ans.8


please tell how?


 

Grade:12

5 Answers

Har Simrat Singh
11 years ago

let x be wt of fe3+ and then 0.96-x will be fe2+ 

the molecule / crystal is neutral therefore net positive = net negative

3x+ 2(0.96-x) = 2

x = 0.08

or 8 percent

Vikas TU
11 years ago

96 Fe atom are associated with 100 O atoms . out of 96 Fe atoms suppose Fe present as Fe+2 = x and Fe present as Fe+3 = 96 -x

total charge on x Fe+2 ions and (96-x) Fe+3 ions should be equal to charge on 100 O-2 ions...this is because overall molecule is neutral

so x X 2 + (96-x) X 3 = 100 X 2

2x + 3(96-x) = 200

2x + 288 - 3x = 200

-x = 200 - 288

x = 88

so out of 96 atoms 88 are Fe+ and 8 are Fe+3...

so fraction of Fe present as Fe+2 = 88/96 X 100 = 91.67 %

and fraction of Fe present as Fe+3 = 8/96 X 100 = 8.33 %

 

 

PLZ  APPROVE!

THNX!

:)

 


Roshan Mohanty
10 years ago

i can write it as Fe96O100
Let Fe3+ be x and Fe2+ be 96-x
by neutralising the charge
3x + 2(96-x) = 200
=> x= 8
and 96-x = 88
so percentage of Fe(3+) is 8.33 %

Please approve 

Kushagra Madhukar
3 years ago
Hello student
As we know, Fe exist in two forms, namely Fe2+ and Fe3+
now the given compound is Fe0.96O
Let us assume 1 mole of Fe0.96O
Therefore moles of Fe = 0.96 ; and moles of O = 1
Let the moles of Fe in Fe3+ form = x
Therefore, the moles of Fe in Fe2+ form = 0.96 -x
 
Now, since net charge on the compound = 0
Hence,
3x + 2(0.96-x) -2 = 0
or, x = 0.08
Hence percentage of Fe3+ ions = 0.08/0.96 * 100
                                                   = 8.33%
 
Hope it helps
Regards,
Kushagra
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your question.
 
As we know, Fe exist in two forms, namely Fe2+ and Fe3+
now the given compound is Fe0.96O
Let us assume 1 mole of Fe0.96O
Therefore moles of Fe = 0.96 ; and moles of O = 1
Let the moles of Fe in Fe3+ form = x
Therefore, the moles of Fe in Fe2+ form = 0.96 -x
 
Now, since net charge on the compound = 0
Hence,
3x + 2(0.96-x) -2 = 0
or, x = 0.08
Hence percentage of Fe3+ ions = 0.08/0.96 * 100
                                                   = 8.33%
 
Hope it helps.
Thanks and regards,
Kushagra